OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#12546 — th2 - gsw
Incident Report for Web Cloud
Resolved
Some connections are no longer mounting between th2 and global switch.

We're going to remove and reinsert some optical cables to see if the connections remount.


Update(s):

Date: 2015-02-05 15:34:07 UTC
The TH2 <> GSW connections are back up.

Date: 2015-02-05 09:24:40 UTC
One of the connections we remounted yesterday is continuing to flap. We therefore have 3 connections to fix at TH2.

Date: 2015-02-04 17:24:46 UTC
Out of the 5 connections that are down, we've remounted 3 from GlobalSwitch. We just need to remount 2 to TH2.

A flap on a connection was able to cause several uplinks/downlinks on the pair 10 between 16:14 and 16:17 (CEST).

We will keep the task open for the TH2 intervention.
Posted Feb 04, 2015 - 15:20 UTC